Lucido Products markets two computer games: Claimjumper and Makeover. A contribution format income statement for a recent month for the two games appears below:

 Sales Variable expenses Claimjumper $ 116,000 42,760 Makeover $ 58,000 9,440 Total $ 174,000 52,200 Contribution margin Fixed expenses $ 73,240 $ 48,560 121,800 77,280 Net operating income $ 44,520 


Required: 

1. Compute the overall contribution margin (CM) ratio for the company. Overall CM ratio 70% 

2. Compute the overall break-even point for the company in dollar sales. (Do not round intermediate culations. Round your final answer to the nearest dollar amount.) Overall break-even point $ 110,400

3. Complete the contribution format income statement at break-even point for the company showing the appropriate levels of sales for the two products. (Do not round intermediate calculations. Round your final answers to the nearest dollar amount.) 
Answer #1

Requirement 1

Contribution margin ratio 70%

Requirement 2

Overall Break-even point $ 110,400.00

Requirement 3

Lucido Products
Contribution income statement
Claimjumper Makeover Total
Sales $          73,600 $              36,800 $        110,400
Variable expenses $          27,130 $                 5,990 $           33,120
Contribution margin $          46,470 $              30,810 $           77,280
Fixed cost $           77,280
Net operating income (loss) $                    -  

Working

Sales $ 174,000.00
Variable expenses $    52,200.00
Contribution margin $ 121,800.00
Contribution margin ratio 70%
Fixed cost $    77,280.00
Overall Break-even point $ 110,400.00

.

Lucido Products
Contribution income statement
Claimjumper Makeover
Sales =116000/174000*110400 =58000/174000*110400
Variable expenses =42760/116000*73600 =9440/58000*36800